OpenStudy (anonymous):

Find the Limit of lim =((1/x)-(1/5))/x-5 X=>5

OpenStudy (anonymous):

\[\lim_{x \rightarrow 5}=(1/x)-(1/5))/x-5\]

OpenStudy (anonymous):

have you covered l'hopitals rule ?

OpenStudy (anonymous):

no not at all, but i figured out what I did wrong, if you don't mind. What exactly is the rule?

OpenStudy (anonymous):

if you havent covered you can solve it other way , just lhopitals rule makes it easier it is : if limit is of form 0/0 or infinity/infinity , you take the derivative of numerator and denominator and the limit wont change but might make it easier to solve

OpenStudy (anonymous):

what did u get for result ?

OpenStudy (jdoe0001):

\(\bf \lim_{x \rightarrow 5} \cfrac{ \frac{1}{x} - \frac{1}{5} }{ x-5 } \implies \cfrac{\frac{5-x}{5x} }{x-5} \implies \cfrac{5-x}{5x} \times \cfrac{1}{x-5}\\ \lim_{x \rightarrow 5}\cfrac{-(x-5)}{5x} \times \cfrac{1}{x-5}\)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 5} \frac{ \frac{ 5-x }{ 5x } }{ x-5 }\rightarrow \frac{ -1 }{ 5x }=\frac{ -1 }{ 25 }\]

OpenStudy (anonymous):

sorry @jdoe0001 I didn't see your reply

OpenStudy (jdoe0001):

np, it just came about 2secs before yours :)